Sabir is going to university to study to become a pharmacist.his father decided to buy him a new Toyota Augo as a reward for his wonderful grade 12 results.The Toyota Augo is currently on special for the reasonable price of R174 900 including VAT.They decided to buy the car on hire-purchase and pay of the balance in 60 montly installment.Simple interest is charged at a rate of 14% p.a. A deposit of 10% is required.what is the Loan amount after the deposit has been paid


1
Expert's answer
2021-03-29T14:14:14-0400

We will find the amount after the deposit is made:

"174 900-174 900\times0.1=157 410"

Find the loan amount:

Monthly payment:

"x=S\times(r+\frac{r}{(1+r)^n-1})"

n=60

S= 157 410

"r=\frac{0.14}{12}=0.0117"

"x=S\times(r+\frac{r}{(1+r)^n-1})=157 410\times(0.0117+\frac{0.0117}{(1+0.0117)^{60}-1})=3662.66"

Find the loan amount:

"3662.66\times60=219 759.32"